two spheres A and B of masses M1 and M2 respectively collide.A is at any rest initially and B is moving with velocity V along x-axis after collision B has a velocity V/2 in a direction perpendicular to the original direction The Mass A moves after the collision in the direction

Grade:12th pass

2 Answers

Arun
6 years ago
Here we will use 2- dimensional momentum conservation
m1* 0 + m2* v = m2*0 + m1*Vxa
Vxa = m2*v/m1
 
as B is moving in y direction only after collision
Similarly
m1*0 + m2*0 = m1* Vya + m2* v/2
So
Vya = - m2*v /2*m1
Hence
Vya/Vxa = -1/2 = tan\theta
So mass A will move tan --¹(-1/2) with positive x axis
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

We will use two dimensional moment conversation.
let as assume the direction B move initially is positive x direction
so we can write,
m1×0+m2×V= m2×0+m1×VxA as B is moving y direction only a after collision
so
VxA = m2/m2 V towards positive direction
and similarly m1 ×0+ m2×0 = m1 × VyA +m2×V/2
VyA = - m2v/ 2m1
VyA/VxA = - 1/2 = tanθ
massA will move than tan^-1 (-0.5) = 26.5° with positive x axis.

Thanks and Regards